OpenStudy (nightmarenight):

Someone heeeeelp? Please? I'll fan and give a medal.... How many solutions does this system of equations have? y = –3x + 7 y = –3x – 6 infinitely many 2 1 0

OpenStudy (superdavesuper):

Hahahaa this is the WRONG but easy way to answer this: Subtract the two equations and u get y - y = -3x - (-3x) +7 -6 so 0 = 0 +7 -6 which is impossible. So the answer is 0 - there is no solution.

OpenStudy (nightmarenight):

Ah! Thank you c: @superdavesuper